12. Verity that:
r + Y1 + r - axyz
= 2 (x + y + z)[(x -y'J' + (y - z>2 + (z - x>31

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Consider identity r3 + y3 + z3 - 3%yz = (x + y + z> t + ez2 + x2 - 2.u)I = 2(x + Y + z) ((x - y>2 + (y - z>2 + (z -x>21.